# Probing Picohertz Gravitational Waves with Pulsars

### v1

@@ I Introduction
Gravitational waves (GW) have opened an entirely new observational window on the Universe. [1, 2, 3] Over the past decade, multiple detection techniques have dramatically advanced the field [4]. Ground-based interferometers such LIGO, Virgo, and KAGRA operate in the Hz to kHz range, successfully detecting hundreds of stellar-mass compact binary coalescences [5]. Space-based interferometry missions, notably LISA [6], are designed to explore the mHz regime. Pulsar timing arrays (PTAs) [7, 8, 9, 10] operate within the nHz regime and have recently uncovered evidence for the stochastic gravitational wave background (GWB) likely produced by supermassive black hole binaries (SMBHB), and potentially other cosmological sources. At the lowest frequencies, between $10^{-18}$ and $10^{-16}$ Hz, the imprints of primordial GWs can be inferred from polarization measurements in the cosmic microwave background (CMB) [11].
This leaves a significant gap in GW frequency band between $10^{-16}$ Hz and 1 nHz, called the picohertz (pHz) regime, where exciting physics may be hiding. First of all, here lies a population of pHz frequency SMBHBs – progenitors of the nHz SMBHB population found in the PTA band [12]. All these binaries emit continuous GWs (CW), and a GWB arises from the superposition of all these individual CW sources. In the pHz regime, environmental effects likely dominate over GW emission in the binary evolution, which should influence the SMBHB demographics. Detecting GWs in this pHz frequency regime can teach us about the SMBHB population, and in turn give us insight into the physics driving SMBHB evolution.
A range of beyond-the-Standard-Model and early-Universe processes could also contribute to the GW signals in this band. These include networks of cosmic strings [13, 14], first-order phase transitions associated with symmetry breaking events in the early Universe [15, 16], turbulent processes during the quantum chromodynamics epoch [17, 18, 19], and primordial GWs generated during inflation [20, 21] and subsequent preheating phases [22]. Detection of pHz GWs would thus open a new observational frontier for testing new physics.
@@
The extensive data accumulated by PTA campaigns, combined with increasingly sophisticated pulsar noise models [23, 24, 25], provide an unprecedented opportunity to explore the pHz frequency regime. The concept that pHz GWs can induce measurable perturbations in pulsar timing parameters was first proposed by Bertotti et al. [26], and later expanded upon by DeRocco and Dror [27, 28] in a modern theoretical framework. Here we present a new Bayesian framework that uses MSP timing parameters to search for CWs in the pHz frequency band. We develop an analytic model that describes the CW-induced drifts in the observed pulsar parameters $\dot{P}_{b}$ and $\ddot{P}$.
For our study, we extend an existing GWB model for SMBHBs [29] to pHz frequencies. The population is matched in a self-consistent way to the evidence for a GWB in the NANOGrav 15-year data [10]. From this extended model, we derive the characteristic strain of the background and generate synthetic realizations of pHz CW sources. Our analysis improves the upper limits on CW strain amplitudes by up to an order of magnitude relative to DeRocco and Dror [28], and achieves a new leading constraint. Furthermore, through detailed simulations, we explore the CW scaling laws in the pHz regime and forecast strain amplitude sensitivities accessible by next-generation telescopes such as the Square Kilometre Array (SKA). We show that by timing new pulsars and improving timing precision with SKA, the technique presented here could potentially detect individual pHz sources, opening a new avenue to probe fundamental physics not accessible by any other means.
@@ II.1 Analytical Model
To leading order, the observed derivative of the binary orbital period can be written as [31]
@@
In principle, there are kinetic and galactic contributions to $\ddot{P}_{\mathrm{obs}}$ as well, but the non-relativistic nature of the galaxy makes them subdominant to $j_{\mathrm{GW}}$ by several orders of magnitude [28]. Intrinsic contributions due to magnetic dipole braking are of order $(\dot{P}/P)^{2}$, which typically evaluate to $\sim 10^{-35}s^{-2}$, much below the observed uncertainty of $\ddot{P}_{\mathrm{obs}}/P$ [32]. The higher order parameters such as $\dddot{P}$ can be derived in a similar fashion, and could be used for GW search in the future analysis. Explicit expressions for $a_{\mathrm{GW}}$ and $j_{\mathrm{GW}}$ induced by a CW are derived in Appendix A.
@@ II.2 Bayesian Analysis
For this study, we consider GWs emitted by SMBHBs in a circular orbit. However, the method is easily generalizable to other CW sources. We have used 7 parameters characterizing the CW signal: $h_{0}$, the GW strain amplitude; $f_{\text{GW}}$, the GW frequency; and $\mathbb{\Theta}$, which consist of $\{l,b,i,\psi,\Phi_{0}\}$. In order $\mathbb{\Theta}$ include the galactic longitude and latitude $l,b$, the orientation of the GW source $i,\psi$, and the orbital phase of the SMBH binary $\Phi_{0}$. Due to the relatively high dimensionality of the problem, we adopt a Bayesian approach.
We construct a likelihood making the standard assumption that pulsar uncertainties are uncorrelated [32]. The likelihood is given by
@@
$P(\theta)=\pi(\alpha)\pi(\delta)\pi(i)\pi(\psi)\pi(\phi_{0})\pi(f)\pi(h)\,.$ (7)
@@
$l$ uniform $0^{\circ}\leq l<360^{\circ}$
$b$ $\frac{1}{2}\sin(b)$ $0^{\circ}\leq b<180^{\circ}$
@@
$\psi$ uniform $\frac{\pi}{2}<\psi<\frac{3\pi}{2}$
$\Phi_{0}$ uniform $0<\Phi_{0}<2\pi$
@@ III Supermassive Black Hole Binary Population Model
In order to understand the detection prospects of our method, we extend the SMBHB population synthesis model from [29] to the pHz regime. This model takes the GWB amplitude measurement as an input, and uses it to constrain SMBHB population characteristics, including the minimum black hole mass contributing to the background, the number density of SMBHBs per unit volume, and the volume that encloses the GWB. Here we use this population synthesis code to anchor the pHz SMBHB population to the nHz population, creating a self-consistent model of our GW sources.
In this model, we use dual active galactic nuclei (AGN) as tracers of SMBHBs: we assume a fixed ratio between dual AGN density $\Phi_{\mathrm{Q}}$ and SMBHB density $\Phi_{\mathrm{BHB}}$, depending on a mass ratio distribution $p(q)$:
$\frac{d^{3}\Phi_{\mathrm{BHB}}}{dM_{\mathrm{BH,1}}\,dz\,dq}=R\frac{d^{2}\Phi_{\mathrm{Q}}}{dM_{\mathrm{BH}}\,dz}\,p(q)\,$, (8)
where $M_{\mathrm{BH,1}}$ is the mass of the primary, $M_{\mathrm{BH}}$ is the total mass of the binary, and $z$ is the redshift, $q$ is the mass ratio. The model parameters [29] including the constant of proportionality $R$ are then extracted by fitting the integrated GWB characteristic strain to the evidence of GWB in the NANOGrav 15-year data.
Towards the pHz regime, several effects that shrink the SMBHB orbits are likely to compete with or even dominate over the GW emission of the binary system. The orbits of the SMBHs which will eventually comprise the binary initially evolve under dynamical friction [33]. As the SMBHs approach a separation of $\sim 100$ pc, their evolution becomes dominated by three-body interactions with individual stars, also known as “stellar hardening”. At separations of $\sim 1$ pc stellar hardening becomes inefficient as the SMBHB clears its vicinity of nearby stars. At this stage interactions between the SMBHB and nearby gas may become important for further shrinking the binary orbit, though the details of this process remain an open question.
In our model, we include both dynamical friction and stellar hardening effects. We do not include the gas hardening effect and leave it for future investigation. Our method of pHz GW detection would be an important step to probe the various models of SMBHB evolution, manifested as a turnover that deviates from the power-law spectrum of GWB in the PTA frequency band.
@@ IV Dataset
For the $\dot{P}_{b}$ analysis, we use the updated data from [34], which includes 29 pulsars. For our purposes, we independently estimate the Galactic accelerations of the pulsars using the MWPotential2014 model as implemented in the galpy package [35].
@@
The original data set from [36] used in DeRocco and Dror [28] contains 14 pulsars, 13 of which are also included in our updated sample. After cross-checking with other recent data releases, we adopt the more updated measurements from [34] for all overlapping pulsars. In total, we obtain 30 pulsars with updated timing measurements, which is double the number used in the analysis of DeRocco and Dror [28].
@@
Figure 2: CW strain amplitude upper limits and forecast detection sensitivities of pulsar timing parameters $\dot{P}_{b}$ and $\ddot{P}$. The fiducial CW sources are modeled as circular orbit SMBHBs. The current upper limit in this work is an order of magnitude better than best constraints from previous work in this frequency range [28]. Forecast sensitivities with SKA data provide another order of magnitude improvement by increase of data size alone, without accounting for the improvement in data uncertainty. Current results are comparable to PTA upper limit on CWs (purple dashed line) for certain frequencies. Expected individual CW sources (denoted by the stars) are at least 3 orders of magnitude in strain amplitude below the current constraint according to our SMBHB population model. The orange solid line is the GWB spectrum drawn from our population model by matching self-consistently with GWB at $f_{\mathrm{GW}}=1~\mathrm{yr}^{-1}$ detected by NANOGrav. The shaded regions centered on this curve denotes the uncertainty quantiles. The black dashed line represents the power-law GWB spectrum matched to NANOGrav GWB at $f_{\mathrm{GW}}=1~\mathrm{yr}^{-1}$.
@@
Pulsar distance measurements are very important in this analysis. In this work, we require $fd_{p}\leq 1$, or that we constrain the pulsar distance to less than a radiation wavelength, which helps mitigate the degeneracies associated with $2\pi$ periodicity of trig functions. For the pHz regime, this is typically $\sim 100$ pc. About half of the pulsars here meet this criterion and are used in this analysis, as we use $f_{\mathrm{GW}}=100$ pHz to define the cutoff. When we perform our MCMC analysis, we use the pulsar distance measurements as a prior, and sample in the restricted error range in the analysis.
@@
Additionally, to ensure that the Taylor expansions used in the analytic modeling remain valid, we cut off the analysis at $(4T_{\mathrm{max}})^{-1}$, consistent with [28], where $T_{\text{max}}$ is the longest observation time in the data.
@@
We account for the Shklovskii contribution $a_{\text{kin}}$ using the analytic formula in Section II. The uncertainty in $d$ dominates the uncertainty in $a_{\text{kin}}$, allowing us to treat $\mu$ as effectively fixed with a small uncertainty. When sampling $d$, we will propagate the proper motion uncertainty and add it in quadrature to the total uncertainty in $a_{\text{kin}}$. We verify that our cut on $\dot{P}_{b}$ data has resulted in a better upper limit by simulating null distributions, finding that including all pulsars shifts the null Bayes factor distribution towards larger values, which in turn biases the posterior toward higher strain amplitudes. The values of $\mu$ used in our analysis are taken from [34].
For the $\ddot{P}$ analysis, we use the same data as in [28], given by [37] and [38]. However, since the distances of most pulsars in this set are estimated using dispersion measures and models of the Galactic free electron density, they lack well-defined uncertainties [39]. For all these pulsars, we conservatively assume a $40\%$ relative uncertainty on the distance.
@@
We also include an estimate of the red noise following the procedure described in [28]. The red noise is treated as a stochastic background given by a power-law spectrum, and we estimate the ensemble variance of the two timing parameters of interest due to the red noise. For the current data, red noise contributions are negligible in the $\dot{P}_{b}$ measurements but relevant for $\ddot{P}$. We incorporate the red noise contribution into the total uncertainty budget for $\ddot{P}$. We emphasize that as timing precision improves, red noise may no longer remain subdominant to other sources of uncertainty in $\dot{P}_{b}$ and will need to be treated carefully in future analyses. The full data can be found in Appendix C.

### v2

@@ I Introduction
Gravitational waves (GW) have opened an entirely new observational window on the Universe. [maggiore2008gravitational, afzal2023nanograv, 2016gravitational] Over the past decade, multiple detection techniques have dramatically advanced the field [sathyaprakash2021gravitational]. Ground-based interferometers such LIGO, Virgo, and KAGRA operate in the Hz to kHz range, successfully detecting hundreds of stellar-mass compact binary coalescences [Abadie_2010]. Space-based interferometry missions, notably LISA [Amaro-Seoane2017], are designed to explore the mHz regime. Pulsar timing arrays (PTAs) [Antoniadis2022, Chen2021, Goncharov2021, nanograv] operate within the nHz regime and have recently uncovered evidence for the stochastic gravitational wave background (GWB) likely produced by supermassive black hole binaries (SMBHB), and potentially other cosmological sources. At the lowest frequencies, between $10^{-18}$ and $10^{-16}$ Hz, the imprints of primordial GWs can be inferred from polarization measurements in the cosmic microwave background (CMB) [Planck2018Inflation].
This leaves a significant gap in GW frequency band between $10^{-16}$ Hz and 1 nHz, called the picohertz (pHz) regime, where exciting physics may be hiding. First of all, here lies a population of pHz frequency SMBHBs – progenitors of the nHz SMBHB population found in the PTA band [Begelman1980]. All these binaries emit continuous GWs (CW), and a GWB arises from the superposition of all these individual CW sources. In the pHz regime, environmental effects likely dominate over GW emission in the binary evolution, which should influence the SMBHB demographics. Detecting GWs in this pHz frequency regime can teach us about the SMBHB population, and in turn give us insight into the physics driving SMBHB evolution.
A range of beyond-the-Standard-Model and early-Universe processes could also contribute to the GW signals in this band. These include networks of cosmic strings [Ghoshal2023, Chang2020], first-order phase transitions associated with symmetry breaking events in the early Universe [Freese2023, Weir2018], turbulent processes during the quantum chromodynamics epoch [Neronov2021, Brandenburg2021, Moore2021], and primordial GWs generated during inflation [Liddle2000, Kamionkowski2016] and subsequent preheating phases [Boyle2005]. Detection of pHz GWs would thus open a new observational frontier for testing new physics.
@@
The extensive data accumulated by PTA campaigns, combined with increasingly sophisticated pulsar noise models [Goncharov2024, Larsen2024, Agazie2023], provide an unprecedented opportunity to explore the pHz frequency regime. The concept that pHz GWs can induce measurable perturbations in pulsar timing parameters was first proposed by Bertotti1983, and later expanded upon by DeRocco2023, DeRoccoDror2024 in a modern theoretical framework. Here we present a new Bayesian framework that uses MSP timing parameters to search for CWs in the pHz frequency band. We develop an analytic model that describes the CW-induced drifts in the observed pulsar parameters $\dot{P}_{b}$ and $\ddot{P}$.
For our study, we extend an existing GWB model for SMBHBs [Casey-Clyde2022] to pHz frequencies. The population is matched in a self-consistent way to the evidence for a GWB in the NANOGrav 15-year data [nanograv]. From this extended model, we derive the characteristic strain of the background and generate synthetic realizations of pHz CW sources. Our analysis improves the upper limits on CW strain amplitudes by up to an order of magnitude relative to DeRoccoDror2024, and achieves a new leading constraint. Furthermore, through detailed simulations, we explore the CW scaling laws in the pHz regime and forecast strain amplitude sensitivities accessible by next-generation telescopes such as the Square Kilometre Array (SKA). We show that by timing new pulsars and improving timing precision with SKA, the technique presented here could potentially detect individual pHz sources, opening a new avenue to probe fundamental physics not accessible by any other means.
@@ II.1 Analytical Model
To leading order, the observed derivative of the binary orbital period can be written as [Bell1996]
@@
In principle, there are kinetic and galactic contributions to $\ddot{P}_{\mathrm{obs}}$ as well, but the non-relativistic nature of the galaxy makes them subdominant to $j_{\mathrm{GW}}$ by several orders of magnitude [DeRoccoDror2024]. Intrinsic contributions due to magnetic dipole braking are of order $(\dot{P}/P)^{2}$, which typically evaluate to $\sim 10^{-35}s^{-2}$, much below the observed uncertainty of $\ddot{P}_{\mathrm{obs}}/P$ [LorimerKramer2012]. The higher order parameters such as $\dddot{P}$ can be derived in a similar fashion, and could be used for GW search in the future analysis. Explicit expressions for $a_{\mathrm{GW}}$ and $j_{\mathrm{GW}}$ induced by a CW are derived in Appendix A.
@@ II.2 Bayesian Analysis
For this study, we consider GWs emitted by SMBHBs in a circular orbit. However, the method is easily generalizable to other CW sources. We have used 7 parameters characterizing the CW signal: $h_{0}$, the GW strain amplitude; $f_{\text{GW}}$, the GW frequency; and $\mathbb{\Theta}$, which consist of $\{\alpha,\delta,i,\phi,\Psi_{0}\}$. In order $\mathbb{\Theta}$ include the right ascension and declination $\alpha,\delta$, the orientation of the GW source $i,\phi$, and the orbital phase of the SMBH binary $\Psi_{0}$. Due to the relatively high dimensionality of the problem, we adopt a Bayesian approach.
We construct a likelihood making the standard assumption that pulsar uncertainties are uncorrelated [LorimerKramer2012]. The likelihood is given by
@@
$P(\theta)=\pi(\alpha)\pi(\delta)\pi(i)\pi(\phi)\pi(\Psi_{0})\pi(f_{\text{GW}})\pi(h_{0})\,.$ (7)
@@
$\alpha$ uniform $0^{\circ}\leq\alpha<360^{\circ}$
$\delta$ $\frac{1}{2}\cos(b)$ $90^{\circ}\leq\delta<90^{\circ}$
@@
$\phi$ uniform $\frac{\pi}{2}<\phi<\frac{3\pi}{2}$
$\Psi_{0}$ uniform $0<\Psi_{0}<2\pi$
@@ III Supermassive Black Hole Binary Population Model
In order to understand the detection prospects of our method, we extend the SMBHB population synthesis model from [Casey-Clyde2022] to the pHz regime. This model takes the GWB amplitude measurement as an input, and uses it to constrain SMBHB population characteristics, including the minimum black hole mass contributing to the background, the number density of SMBHBs per unit volume, and the volume that encloses the GWB. Here we use this population synthesis code to anchor the pHz SMBHB population to the nHz population, creating a self-consistent model of our GW sources.
In this model, we use dual active galactic nuclei (AGN) as tracers of SMBHBs: we assume a fixed ratio between dual AGN density $\Phi_{\mathrm{Q}}$ and SMBHB density $\Phi_{\mathrm{SMBHB}}$, depending on a mass ratio distribution $p(q)$:
$\frac{d^{3}\Phi_{\mathrm{SMBHB}}}{dM_{\mathrm{BH,1}}\,dz\,dq}=R\frac{d^{2}\Phi_{\mathrm{Q}}}{dM_{\mathrm{BH}}\,dz}\,p(q)\,$, (8)
where $M_{\mathrm{BH,1}}$ is the mass of the primary, $M_{\mathrm{BH}}$ is the total mass of the binary, $z$ is the redshift, and $q$ is the mass ratio. The model parameters [Casey-Clyde2022] including the constant of proportionality $R$ are then extracted by fitting the integrated GWB characteristic strain to the evidence of GWB in the NANOGrav 15-year data.
Towards the pHz regime, several effects that shrink the SMBHB orbits are likely to compete with or even dominate over the GW emission of the binary system. The dual–AGN that contain the SMBHs which will eventually comprise the binary initially evolves under dynamical friction [BinneyTremaine2008]. As the SMBHs approach a separation of $\sim 100$ pc, the evolution of the dual–AGN becomes dominated by three-body interactions with individual stars, also known as “stellar hardening”. At separations of $\sim 1$ pc stellar hardening becomes inefficient as the SMBHs clears their vicinity of nearby stars. At this stage interactions between the SMBHB and nearby gas may become important for further shrinking the SMBH separation, though the details of this process remain an open question. Hard binaries of SMBHs start to form at this separation. For a SMBHB of mass $10^{8}\sim 10^{10}\mathrm{{M}}_{{\odot}}$ and separation $\sim 1~\text{pc}$, the emitted GW frequencies are $\sim 10-100~\text{pHz}$.
In our model, we include both dynamical friction and stellar hardening effects. We do not include the gas hardening effect and leave it for future investigation. Our method of pHz GW detection would be an important step to probe the various models of SMBHB evolution, manifested as a turnover that deviates from the power-law spectrum of GWB in the PTA frequency band. The GWB from our population model is orders of magnitude below the CW upper limits, so the noise due to GWB can be reasonably ignore in this work.
@@ IV Dataset
For the $\dot{P}_{b}$ analysis, we use the updated data from [Moran:2023], which includes 29 pulsars. For our purposes, we independently estimate the Galactic accelerations of the pulsars using the MWPotential2014 model as implemented in the galpy package [Bovy2015].
@@
The original data set from [Chakrabarti_2021] used in DeRoccoDror2024 contains 14 pulsars, 13 of which are also included in our updated sample. After cross-checking with other recent data releases, we adopt the more updated measurements from [Moran:2023] for all overlapping pulsars. In total, we obtain 30 pulsars with updated timing measurements, which is double the number used in the analysis of DeRoccoDror2024. The observed $\dot{P}_{b}$’s range in order of magnitude from $10^{-21}$ to $10^{-8}$, with the majority being around $10^{-14}\sim 10^{-12}$. We subtract non-GW contributions from the observed $\dot{P}_{b}$ and report the GW induced $\Delta\dot{P}_{b}/P_{b}$ data immediately used for analysis. The timing baselines of the pulsars range from 4 years to 22 years, with $\mu s$ level RMS timing residuals. For a full list of the pulsars, see Table 2.
@@
Figure 2: CW strain amplitude upper limits and forecast detection sensitivities of pulsar timing parameters $\dot{P}_{b}$ and $\ddot{P}$. The fiducial CW sources are modeled as circular orbit SMBHBs. The current upper limit in this work is an order of magnitude better than best constraints from previous work in this frequency range [DeRoccoDror2024]. Forecast sensitivities with SKA data provide another order of magnitude improvement by increase of data size alone, without accounting for the improvement in data uncertainty. Current results are comparable to PTA upper limit on CWs (purple dashed line) for certain frequencies. Expected individual CW sources (denoted by the stars) are at least 3 orders of magnitude in strain amplitude below the current constraint according to our SMBHB population model. The orange solid line is the GWB spectrum drawn from our population model by matching self-consistently with GWB at $f_{\mathrm{GW}}=1~\mathrm{yr}^{-1}$ detected by NANOGrav. The shaded regions centered on this curve denotes the uncertainty quantiles. The black dashed line represents the power-law GWB spectrum matched to NANOGrav GWB at $f_{\mathrm{GW}}=1~\mathrm{yr}^{-1}$.
@@
Pulsar distance measurements are essential for reducing degeneracies in this analysis. If the resulting dataset is large enough, we require $f\Delta d_{p}\leq 1$; namely, we constrain the pulsar distance up to a precision less than a radiation wavelength, which helps mitigate the degeneracies associated with $2\pi$ periodicity of trig functions. For the pHz regime, this typically corresponds to a wavelength of $\sim 100$ pc. Half of the pulsars in the $\dot{P}_{b}$ data meet this criterion and are used in the analysis, as we use $f_{\mathrm{GW}}=100$ pHz to define the cutoff. When we perform our MCMC analysis, we use the pulsar distance measurements as a prior, and sample in the restricted error range in the analysis.
@@
Additionally, to ensure that the Taylor expansions used in the analytic modeling remain valid, we cut off the analysis at $(4T_{\mathrm{max}})^{-1}$, consistent with [DeRoccoDror2024], where $T_{\text{max}}$ is the longest observation time in the data.
@@
We account for the Shklovskii contribution $a_{\text{kin}}$ using the analytic formula in Section II. The uncertainty in $d$ dominates the uncertainty in $a_{\text{kin}}$, allowing us to treat $\mu$ as effectively fixed with a small uncertainty. When sampling $d$, we will propagate the proper motion uncertainty and add it in quadrature to the total uncertainty in $a_{\text{kin}}$. We verify that our cut on $\dot{P}_{b}$ data has resulted in a better upper limit by simulating null distributions, finding that including all pulsars shifts the null Bayes factor distribution towards larger values, which in turn biases the posterior toward higher strain amplitudes. The values of $\mu$ used in our analysis are taken from [Moran:2023].
For the $\ddot{P}$ analysis, we use the same data as in [DeRoccoDror2024], given by [Desvignes2016] and [Reardon2016]. A full list can be found in Table 3.The pulsars have timing baselines ranging from 7 to 18 years, with a typical RMS timing residual of a few $\mu s$. We use the parallax measurement of the pulsar distance whenever it is available. Otherwise, we estimate using dispersion measure (DM) and a model of Galactic free electron density model. A total of 13 pulsar distances in the data are estimated using DM and the Galactic free electron density model YMW 17 [yao2017new]. Since they lack a well-defined uncertainty in distance, we conservatively assume a $40\%$ relative uncertainty on these distances. Only one fourth of the pulsar distances in the $\ddot{P}$ data satisfy $f\Delta d_{p}\leq 1$, and dropping the majority of the data will make the set prone to outliers and lose information from a larger data set. Thus, we retain all pulsars in the $\ddot{P}$ data for analysis.
@@
We also include an estimate of the red noise following the procedure described in [DeRoccoDror2024]. The red noise is treated as a stochastic background given by a power-law spectrum, and we estimate the ensemble variance of the two timing parameters of interest due to the red noise. For the current data, red noise contributions are negligible in the $\dot{P}_{b}$ measurements but relevant for $\ddot{P}$. We incorporate the red noise contribution into the total uncertainty budget for $\ddot{P}$. We emphasize that as timing precision improves, red noise may no longer remain subdominant to other sources of uncertainty in $\dot{P}_{b}$ and will need to be treated carefully in future analyses. The full data can be found in Appendix C.